Solve.
A triangle has sides 3x + 7, 4x - 9, and 5x + 6. Find the polynomial
that represents its perimeter.

Answer by SHUgrad05(58)

Let a=3x + 7; b=4x - 9; c=5x + 6. Find the polynomial
that represents its perimeter.
The perimeter of a triangle is: P=a+b+c
P= 3x+7+4x-9+5x+6
Now simplify:
P=12x+4